# KVComm: Enabling Efficient LLM Communication through Selective KV Sharing

### v1

@@ B.1 Dataset
We provide sample prompts and expected answers for the Countries and Tipsheets datasets in Table 3, which are inspired by Ramesh & Li (2025). We also provide the statistics of all datasets used in our experiments in Table 4. HotpotQA, QASPER, MuSiQuest, and TMATH datasets are randomly sampled from their original datasets to reduce the evaluation cost.
Table 3: Sample prompts and expected answers for Countries and Tipsheets datasets inspired by Ramesh & Li (2025).
@@ B.4 Compared Method Descriptions
  • Natural Language Debate (NLD) (Du et al., 2023): Each model generates an initial answer, and then they iteratively critique each other’s answers in natural language for a fixed number of rounds. Finally, one model produces the final answer based on the entire debate history. Compared to the original setting, we explicitly tell $\mathcal{M}_{s}$ that it has to summarize the context $C$ in its initial answer. We set the number of debate rounds to 2.
@@
  • AC (Ramesh & Li, 2025): Communicate with the last token’s hidden state. Replace the last token’s hidden state of $\mathcal{M}_{r}$ with that of $\mathcal{M}_{s}$. We also test with mean and sum operations.
@@ Appendix D Utilizing All Tokens
Appendix E More Communication Results
@@
Table 6: More communication results of different methods. Best results are bolded, second best underlined (excluding Baseline and Skyline). We report $\mathcal{M}_{r}$ for Baseline and Skyline for fairness. KVComm (0.3/0.5/0.7) denotes selecting 30%/50%/70% of layers’ KV pairs for communication, i.e., $M=\lceil 0.3L\rceil$, $M=\lceil 0.5L\rceil$, $M=\lceil 0.7L\rceil$.
Method Countries Tipsheets HotpotQA QASPER MuSiQuest
MultiField
-QA-en
2WikiM
-QA
TMATH
@@
Appendix F Ablation Study on Selection Strategy
We conduct more ablation studies on the selection strategy by comparing with random selection and selection based on only attention importance scores. The results are shown in LABEL:tab:more_comparison_with_random, which show similar trends as in Section 4.4.
Method Countries Tipsheets HotpotQA QASPER MuSiQuest
MultiField
-QA-en
2WikiM
-QA
TMATH
Random (0.3) $0.02$ $0.35$ $0.24$ 0.07 $0.04$ $0.07$ $0.12$ $0.35$
KVComm (0.3) 0.51 0.93 0.33 0.07 0.11 0.21 0.29 0.37
Random (0.5) $0.49$ $0.76$ $0.58$ $0.15$ $0.29$ $0.29$ $0.27$ $0.36$
KVComm (0.5) 0.62 0.95 0.60 0.29 0.34 0.50 0.37 0.37
Random (0.7) 0.63 $0.88$ 0.76 0.32 0.49 $0.52$ $0.34$ $0.37$
KVComm (0.7) $0.62$ 0.96 $0.69$ $0.29$ $0.39$ 0.53 0.38 0.38
Random (0.3) $0.00$ $0.09$ $0.00$ $0.00$ $0.00$ $0.06$ $0.01$ 0.31
KVComm (0.3) 0.04 0.26 0.02 0.01 0.01 0.09 0.08 0.31
Random (0.5) $0.12$ $0.32$ $0.06$ $0.00$ $0.03$ $0.15$ $0.04$ 0.33
KVComm (0.5) 0.19 0.88 0.28 0.07 0.12 0.26 0.10 0.33
Random (0.7) $0.16$ $0.76$ $0.14$ $0.03$ $0.02$ $0.20$ $0.04$ 0.34
KVComm (0.7) 0.41 0.89 0.41 0.21 0.25 0.29 0.15 0.34
Random (0.3) $0.35$ $0.36$ $0.23$ $0.06$ $0.07$ $0.14$ $0.24$ 0.31
KVComm (0.3) 0.46 0.69 0.59 0.19 0.40 0.35 0.29 0.32
Random (0.5) $0.23$ $0.42$ $0.27$ $0.09$ $0.08$ $0.15$ $0.28$ $0.31$
KVComm (0.5) 0.40 0.92 0.63 0.25 0.44 0.45 0.34 0.35
Random (0.7) $0.18$ $0.94$ $0.51$ $0.23$ $0.35$ $0.47$ $0.30$ $0.34$
KVComm (0.7) 0.19 0.96 0.55 0.26 0.42 0.51 0.31 0.36
Random (0.3) $0.02$ $0.29$ $0.11$ $0.06$ $0.02$ $0.07$ $0.16$ $0.34$
KVComm (0.3) 0.51 0.48 0.47 0.10 0.20 0.17 0.28 0.35
Random (0.5) $0.28$ $0.44$ $0.30$ $0.06$ $0.06$ $0.06$ $0.19$ $0.35$
KVComm (0.5) 0.55 0.79 0.58 0.24 0.27 0.47 0.35 0.36
Random (0.7) $0.54$ 0.81 $0.62$ $0.21$ 0.30 $0.30$ $0.26$ $0.36$
KVComm (0.7) 0.57 $0.80$ 0.65 0.27 $0.29$ 0.48 0.31 0.37
Random (0.3) $0.00$ $0.34$ $0.05$ $0.00$ $0.00$ $0.08$ $0.10$ 0.30
KVComm (0.3) 0.04 0.31 0.06 0.02 0.01 0.19 0.19 0.32
Random (0.5) $0.00$ $0.32$ $0.10$ $0.02$ $0.02$ $0.10$ $0.16$ 0.32
KVComm (0.5) 0.57 0.92 0.49 0.18 0.20 0.40 0.25 0.32
Random (0.7) $0.41$ $0.71$ $0.28$ $0.04$ $0.04$ $0.21$ $0.17$ $0.32$
KVComm (0.7) 0.56 0.98 0.72 0.29 0.48 0.45 0.35 0.33
Random (0.3) $0.01$ $0.35$ $0.18$ $0.04$ $0.03$ $0.12$ $0.21$ $0.30$
KVComm (0.3) 0.06 0.67 0.41 0.12 0.22 0.41 0.23 0.32
Random (0.5) $0.04$ $0.41$ $0.24$ $0.03$ $0.05$ $0.16$ $0.24$ $0.31$
KVComm (0.5) 0.16 0.94 0.52 0.22 0.33 0.47 0.33 0.32
Random (0.7) $0.19$ $0.95$ $0.51$ $0.20$ $0.29$ $0.42$ $0.26$ 0.32
KVComm (0.7) 0.23 0.96 0.54 0.22 0.32 0.47 0.29 0.32
Random (0.3) $0.00$ $0.34$ $0.06$ $0.00$ $0.01$ $0.13$ $0.03$ $0.34$
KVComm (0.3) 0.12 0.95 0.12 0.05 0.04 0.26 0.19 0.36
Random (0.5) $0.03$ $0.79$ $0.29$ $0.06$ $0.09$ $0.32$ $0.16$ $0.35$
KVComm (0.5) 0.55 0.98 0.38 0.15 0.14 0.43 0.28 0.38
Random (0.7) $0.37$ $0.85$ 0.59 $0.21$ 0.27 $0.47$ 0.33 $0.36$
KVComm (0.7) 0.53 0.97 $0.51$ 0.22 $0.25$ 0.49 0.33 0.38
Appendix G Calibration Set Size
Figure 11: Effect of calibration set size. Calibration set size does not significantly affect the test performance.
We investigate how many samples are needed in the calibration set so that the selection strategy can generalize well to the test set. If a smaller calibration set can achieve good performance on the test set, it would be more practical since it would require less cost to obtain the selected layers. We conduct the experiment on Countries, Tipsheets, and HotpotQA datasets using the Llama-3.2-3B model. As the results in Figure 11 show, we can see that using only one sample in the calibration set can already achieve the same performance as using more samples (up to 128 samples). This suggests that our selection strategy can generalize well to the test set even with a very small calibration set. In all other experiments in the paper, we use one sample in the calibration set.
Appendix H Complexity Analysis Details
We compare the computational complexity of our KVComm framework with the Skyline method and the NLD method. Recall that $L$ is the total number of layers in the model, $M$ is the number of selected layers for communication. We use $d$ to denote the hidden dimension of the model, and $|Q|$ and $|C|$ to denote the number of tokens in the query and context, respectively. Suppose $\mathcal{M}_{r}$ would generate $T$ tokens in total, and the number of generated tokens is the same across different methods. For NLD, $\mathcal{M}_{s}$ and $\mathcal{M}_{r}$ would each generate $T_{s}$ and $T_{r}$ tokens for the initial answer, respectively.
Ignoring the embedding, output layers, and other minor components, the computational complexity of prefilling a sequence of length $N$ with a single decoder layer is $O(Nd^{2}+N^{2}d)$, while the complexity of decoding a single token is $O(d^{2}+(N+i)d)$, where $i$ is the index of the generated token. Therefore, the total computational complexity of $\mathcal{M}_{s}$ to process the context $C$ is $O(L(|C|d^{2}+|C|^{2}d))$.
The total computational complexity of KVComm consists of three parts: (1) the complexity of $\mathcal{M}_{s}$ to process the context $C$, which is $O(L(|C|d^{2}+|C|^{2}d))$, (2) the complexity of $\mathcal{M}_{r}$ to process the query $Q$ with the selected $M$ KV pairs from $\mathcal{M}_{s}$, which is $O(L|Q|d^{2}+M(|C|+|Q|)|Q|d+(L-M)|Q|^{2}d)$, and (3) the complexity of $\mathcal{M}_{r}$ to generate $T$ tokens with the selected $M$ KV pairs from $\mathcal{M}_{s}$, which is $O(T(Ld^{2}+M(|C|+|Q|+T)d+(L-M)(|Q|+T)d))$. Therefore, the total computational complexity of KVComm is:
$T(\text{KVComm})=$ $\;O\!\Big(L\left(|C|+|Q|+T\right)d^{2}\Big)$
$+O\!\Big(\big(L\left(|C|^{2}+|Q|^{2}+T^{2}+T|Q|\right)+CM\left(|Q|+T\right)\big)d\Big)$
The computational complexity of Skyline method consists of two parts: (1) the complexity of prefilling the concatenation of the context $C$ and query $Q$, which is $O(L(|C|+|Q|)d^{2}+L(|C|+|Q|)^{2}d)$, and (2) the complexity of decoding $T$ tokens, which is $O(TL(d^{2}+(|C|+|Q|+T)d))$. Therefore, the total computational complexity of the Skyline method is:
$T(\text{Skyline})=$ $\;O\!\Big(L\big(|C|+|Q|+T\big)d^{2}\Big)$
$+O\!\Big(L\Big((|C|+|Q|)^{2}+T\big(|C|+|Q|+T\big)\Big)d\Big)$
The margin of KVComm over Skyline is:
$T(\text{Skyline})-T(\text{KVComm})=$ $\;O\!\Big(|C|d\big(L(2|Q|+T)-M(|Q|+T)\big)\Big)$
For NLD, the total computational complexity consists of three parts: (1) the complexity of $\mathcal{M}_{s}$ to process the context $C$ and generate $T_{s}$ tokens, which is $O(L(|C|d^{2}+|C|^{2}d)+T_{s}L(d^{2}+(|C|+T_{s})d))$, (2) the complexity of $\mathcal{M}_{r}$ to process the query $Q$ and generate $T_{r}$ tokens, which is $O(L(|Q|d^{2}+|Q|^{2}d)+T_{r}L(d^{2}+(|Q|+T_{r})d))$, and (3) the complexity of $\mathcal{M}_{r}$ to process the entire debate history and generate $T$ tokens, which is $O(L((T_{s}+T_{r}+|Q|)d^{2}+(T_{s}+T_{r}+|Q|)^{2}d)+TL(d^{2}+(T_{s}+T_{r}+|Q|+T)d))$. Therefore, the total computational complexity of NLD is:
$T(\text{NLD})=$ $\;O\!\Bigg(L\Big(|C|+2|Q|+2T_{s}+2T_{r}+T\Big)d^{2}\Bigg)$
$+O\!\Bigg(L\Big(|C|^{2}+T_{s}^{2}+|Q|^{2}+T_{r}^{2}+\big(T_{s}+T_{r}+|Q|\big)^{2}$
$\quad\quad+T\big(T_{s}+T_{r}+T+|Q|\big)+T_{s}|C|+T_{r}|Q|\Big)d\Bigg)$
The margin of KVComm over NLD is:
$T(\text{NLD})-T(\text{KVComm})=$ $\;O\!\Big(L\big(2T_{s}+2T_{r}+|Q|\big)d^{2}\Big)$
$+O\!\Bigg(\Bigg(L\Big(T_{s}^{2}+T_{r}^{2}+\big(T_{s}+T_{r}+|Q|\big)^{2}$
$+T_{s}|C|+T_{r}|Q|+T(T_{s}+T_{r})\Big)-CM\big(|Q|+T\big)\Bigg)d\Bigg)$
Appendix I Using One Chunk of Layers
We conduct the same experiment as in Section 4.3 on the HotpotQA dataset using other model pairs in Table 5. The results are shown in Figure 13. We can see that the results are consistent with the observation in Section 4.3.
(a) Llama-3.2-3B-Instruct as both $\mathcal{M}_{s}$ and $\mathcal{M}_{r}$
(b) Qwen2.5-7B-Instruct as both $\mathcal{M}_{s}$ and $\mathcal{M}_{r}$
(c) Qwen2.5-7B-Instruct-Uncensored as $\mathcal{M}_{s}$ and Bespoke-Stratos-7B as $\mathcal{M}_{r}$
(d) EvolCodeLlama-3.1-8B-Instruct as $\mathcal{M}_{s}$ and ToolACE-2-Llama-3.1-8B as $\mathcal{M}_{r}$
(e) Llama-3.2-3B-Instruct-abliterated as $\mathcal{M}_{s}$ and DeepSeek-R1-Distill-Llama-3B as $\mathcal{M}_{r}$
Figure 13: Experiment results of using one chunk of layers for communication on HotpotQA dataset using different model pairs.

### v2

@@ 7 Conclusion
Acknowledgments
This work was supported by the Knut and Alice Wallenberg Foundation through a Wallenberg Scholar Grant to Prof. Dejan Kostić. This work has been partially supported by Vinnova (the Sweden’s Innovation Agency), the Swedish Research Council (agreement No. 2021-04212), and KTH Digital Futures. We thank the anonymous reviewers for their insightful comments and constructive suggestions. We also thank Nicolae Filat for providing an early version of the code for CIPHER comparison and for helpful suggestions on the illustrations.
Computations were enabled by the Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the National Supercomputer Centre. We also acknowledge the EuroHPC Joint Undertaking for awarding this project access to the EuroHPC supercomputer LEONARDO, hosted by CINECA (Italy) and the LEONARDO consortium through an EuroHPC Development Access call.
Reproducibility Statement
We provide detailed descriptions of the datasets, model pairs, and implementation details in Appendix B. The code and synthetic datasets, Countries and Tipsheets, are uploaded to GitHub to facilitate reproducibility upon the publication of this work.
@@ Appendix B Experimental Setup
In this appendix, we provide more details about the experimental setup, including dataset details, implementation details, fine-tuned model pairs, and descriptions of compared methods.
@@ B.1 Dataset
We provide sample prompts and expected answers for the Countries and Tipsheets datasets in Table 3, which are inspired by Ramesh and Li (2025). We also provide the statistics of all datasets used in our experiments in Table 4. HotpotQA, QASPER, MuSiQuest, and TMATH datasets are randomly sampled from their original datasets to reduce the evaluation cost. Extended results on the full datasets are provided in Appendix E.
Table 3: Sample prompts and expected answers for Countries and Tipsheets datasets inspired by Ramesh and Li (2025).
@@ B.3 Model Pairs
9 arcee-ai/Llama-3.1-SuperNova-Lite deepseek-ai/DeepSeek-R1-Distill-Llama-8B Fine-tuned on 1
@@ B.4 Compared Method Descriptions
  • Natural Language Debate (NLD) (Du et al., 2023): Each model generates an initial answer, and then they iteratively critique each other’s answers in natural language for a fixed number of rounds. Finally, one model produces the final answer based on the entire debate history. Compared to the original debate style setting, we use an information-transfer style, which explicitly prompts $\mathcal{M}_{s}$ that it has to summarize the context $C$ in its initial answer. We set the number of debate rounds to 2.
@@
  • AC (Ramesh and Li, 2025): Communicate with the last token’s hidden state. Replace the last token’s hidden state of $\mathcal{M}_{r}$ with that of $\mathcal{M}_{s}$. We also test with mean and sum operations.
@@ Appendix C Token Importance at Different Positions
We add more details and experiments related to Section 2.2.1 in this appendix.
C.1 Detailed Experiment Procedure
We provide a detailed description of the experiment procedure in Section 2.2.1. Considering a model $\mathcal{M}$ with $L$ layers, given an input $X$ with $N$ tokens, we run a partial forward pass until layer $l$ to obtain the hidden states $\{\mathbf{h}_{i}^{l}\}_{i=1}^{N}$. Then, given a specific token position $k$, if we perform the Retain operation, we create a modified set of hidden states $\{\tilde{\mathbf{h}}_{i}^{l}\}_{i=1}^{N}$ as follows:
$\tilde{\mathbf{h}}_{i}^{l}=\begin{cases}\mathbf{h}_{i}^{l},&\text{if }i=k\\ \mathbf{0},&\text{otherwise}\end{cases}$
If we perform the Remove operation, we create the modified set of hidden states $\{\tilde{\mathbf{h}}_{i}^{l}\}_{i=1}^{N}$ as follows:
$\tilde{\mathbf{h}}_{i}^{l}=\begin{cases}\mathbf{0},&\text{if }i=k\\ \mathbf{h}_{i}^{l},&\text{otherwise}\end{cases}$
We then continue the forward pass from layer $l+1$ to layer $L$ using the modified hidden states $\{\tilde{\mathbf{h}}_{i}^{l}\}_{i=1}^{N}$ as input, and obtain the final output of the model. We evaluate the model’s performance on the task with different token positions $k$ and layer $l$.
C.2 More Experiments on Token Importance
@@ Appendix D Utilizing All Tokens
We add more details and experiments related to Section 2.2.2 in this appendix.
D.1 Detailed Experiment Procedure
We provide a detailed description of the experiment procedure in Section 2.2.2. Considering two models $\mathcal{M}_{s}$ and $\mathcal{M}_{r}$, each with $L$ layers, given $C$ and $Q$ as input, we run a partial forward pass of $\mathcal{M}_{s}$ until layer $k$ to obtain the hidden state $\mathbf{H}_{k}^{s}\in\mathbb{R}^{|C|\times d}$ for all tokens in $C$, where $|C|$ is the number of tokens in $C$, and $d$ is the hidden dimension. Another partial forward pass of $\mathcal{M}_{r}$ is run until layer $j$ to obtain the hidden state $\mathbf{H}_{j}^{r}\in\mathbb{R}^{|Q|\times d}$ for all tokens in $Q$, where $|Q|$ is the number of tokens in $Q$. We then modify the hidden states of $\mathcal{M}_{r}$ at layer $j$ by prepending the hidden states from $\mathcal{M}_{s}$ at layer $k$ as follows:
$\tilde{\mathbf{H}}_{j}^{r}=\begin{bmatrix}\mathbf{H}_{k}^{s}\\ \mathbf{H}_{j}^{r}\end{bmatrix}$
We continue the forward pass from layer $j+1$ to layer $L$ using the modified hidden states $\tilde{\mathbf{H}}_{j}^{r}$ as input, and obtain the final output of the model. We evaluate the model’s performance on the task with different layers $k$ and $j$.
D.2 More Experiments on Utilizing All Tokens
@@
Appendix E Experiment results with extended datasets
To further validate the effectiveness of our KVComm framework, we conduct experiments on the full datasets of HotpotQA, QASPER, MuSiQuest, mainly HotpotQA-E, QASPER-E, and MuSiQuest-E. Moreover, we include a new human-created summarization dataset, SAMSum, which represents a different task type. The statistics of these datasets are shown in Table 7. We report the results on these extended datasets in Table 6. The results show similar trends as in Section 4.2, demonstrating the robustness of our KVComm framework across different datasets and tasks.
Table 6: Communication results on extended communication tasks. The best results in each block are in bold, and the second best results are underlined.
Method HotpotQA-E QASPER-E MuSiQuest-E SAMSum
$\mathcal{M}_{s}$: huihui-ai/Llama-3.2-3B-Instruct-abliterated;
$\mathcal{M}_{r}$: suayptalha/DeepSeek-R1-Distill-Llama-3B
Baseline $0.22$ $0.03$ $0.06$ $0.26$
Skyline $0.77$ $0.52$ $0.25$ $0.33$
NLD $0.45$ $0.10$ $0.18$ 0.28
CIPHER $0.51$ $0.10$ $0.20$ 0.28
AC (mean) $0.24$ $0.03$ $0.06$ $0.26$
AC (replace) $0.06$ $0.00$ $0.01$ $0.26$
AC (sum) $0.23$ $0.03$ $0.06$ $0.26$
KVComm (0.3) $0.44$ $0.25$ $0.11$ $0.25$
KVComm (0.5) 0.61 0.36 0.25 0.28
KVComm (0.7) 0.71 0.38 0.30 0.29
$\mathcal{M}_{s}$: Orion-zhen/Qwen2.5-7B-Instruct-Uncensored;
$\mathcal{M}_{r}$: bespokelabs/Bespoke-Stratos-7B
Baseline $0.15$ $0.04$ $0.06$ $0.25$
Skyline $0.58$ $0.27$ $0.10$ $0.35$
NLD 0.24 $0.02$ $0.07$ $0.28$
CIPHER $0.04$ $0.01$ $0.02$ 0.37
AC (mean) $0.03$ $0.00$ $0.00$ $0.01$
AC (replace) $0.00$ $0.00$ $0.00$ $0.00$
AC (sum) $0.03$ $0.00$ $0.00$ $0.04$
KVComm (0.3) $0.02$ $0.00$ $0.01$ $0.18$
KVComm (0.5) $0.21$ 0.14 0.08 $0.30$
KVComm (0.7) 0.40 0.34 0.21 0.35
$\mathcal{M}_{s}$: Orion-ehristoforu/falcon3-ultraset;
$\mathcal{M}_{r}$: huihui-ai/Falcon3-7B-Instruct-abliterated
Baseline $0.21$ $0.06$ $0.06$ $0.27$
Skyline $0.78$ $0.60$ $0.33$ $0.36$
NLD 0.52 $0.10$ $0.28$ $0.28$
CIPHER $0.28$ $0.03$ $0.09$ $0.17$
AC (mean) $0.24$ $0.06$ $0.07$ $0.26$
AC (replace) $0.12$ $0.02$ $0.01$ $0.26$
AC (sum) $0.23$ $0.05$ $0.06$ $0.26$
KVComm (0.3) 0.59 $0.15$ $0.40$ $0.28$
KVComm (0.5) 0.59 0.22 0.46 0.31
KVComm (0.7) 0.59 0.26 0.36 0.32
Table 7: Statistics of extended datasets.
Dataset Size
HotpotQA-E (Yang et al., 2018) 7,405
QASPER-E (Dasigi et al., 2021) 1,726
MuSiQuest-E (Trivedi et al., 2022) 2,417
SAMSum (Gliwa et al., 2019) 819
Appendix F More Communication Results
@@
$\mathcal{M}_{s}$: arcee-ai/Llama-3.1-SuperNova-Lite; $\mathcal{M}_{r}$: deepseek-ai/DeepSeek-R1-Distill-Llama-8B
Baseline $0.07$ $0.30$ $0.11$ $0.01$ $0.03$ $0.09$ $0.16$ $0.23$
Skyline $0.55$ $0.80$ $0.52$ $0.17$ $0.40$ $0.41$ $0.16$ $0.29$
NLD $0.30$ $0.39$ $0.20$ $0.02$ $0.06$ $0.08$ 0.19 $0.22$
CIPHER 0.47 0.71 $0.27$ $0.03$ $0.11$ $0.14$ $0.14$ $0.18$
AC (mean) $0.00$ $0.31$ $0.08$ $0.02$ $0.02$ $0.09$ $0.14$ $0.25$
AC (replace) $0.00$ $0.39$ $0.04$ $0.01$ $0.00$ $0.16$ 0.16 0.28
AC (sum) $0.02$ $0.34$ $0.07$ $0.02$ $0.02$ $0.08$ 0.16 $0.24$
KVComm (0.3) $0.09$ $0.52$ $0.10$ $0.01$ $0.03$ $0.09$ $0.08$ 0.28
KVComm (0.5) $0.41$ 0.76 0.33 0.05 0.21 0.23 $0.09$ 0.29
KVComm (0.7) 0.53 0.76 0.47 0.12 0.28 0.31 $0.14$ 0.29